# Green Federated Learning: A new era of Green Aware AI

### v1

@@ 6. Energy and Carbon Footprint Study in Federated Learning
In the previous section 3, several FL-related challenges and solutions are discussed elaborately. However, how FL can affect the ecosystem is still unknown and unclear. There is a relation between FL design and carbon emission (Qiu et al., 2021). According to the survey, an energy-efficient FL framework reduces the carbon footprint. The potential for ML systems to greatly increase carbon emissions exists. Recent research (Schwartz et al., 2019; Strubell et al., 2020) has recommended a number of mitigating techniques and used case studies to illustrate these potential implications. To more accurately assess the overall energy and carbon footprints of machine learning—in both research and production settings—systematic and precise measurements are required. Among other advantages, accurate accounting of carbon and energy impacts promotes awareness, motivates mitigation efforts, and aligns incentives with energy efficiency (Schwartz et al., 2019). However, energy or carbon emissions data are not routinely reported in the majority of ML research articles. Recent research primarily looks into how machine learning research affects the climate. By analyzing the anticipated power consumption and carbon emissions for a series of case studies, Strubell et al. (Strubell et al., 2020) illustrated the problem of carbon and energy consequences of big NLP models. According to the authors, “researchers should prioritize computationally efficient hardware and algorithms”, “academic researchers need equitable access to computation resources”, and “authors should report training time and sensitivity to hyperparameters.” Similar recommendations are made by Schwartz et al. (Schwartz et al., 2019), who propose floating point operations (FPOs) as a guiding efficiency parameter. A website for calculating carbon emissions based on GPU kind, experiment duration, and cloud provider was just made available by Lacoste et al. (Lacoste et al., 2019). Similarly, FL - a cooperative machine learning method that uses data from decentralized entities to train a centralized model—can also be resource-intensive and have a significant carbon impact, mainly when used on a large scale. In contrast to centralized artificial intelligence (AI), which may reliably utilize renewable energy sources at carefully positioned data centers, cross-device FL can utilize up to hundreds of millions of end-user devices spread throughout the globe that use a variety of energy sources. According to a recent study, training a model with FL can emit up to 80 kg of carbon dioxide equivalent (CO2e), which is more than developing a larger transformer model in a centralized training setup using AI accelerators (Wu et al., 2022b). Several reasons contribute to the relative inefficiency: the overhead of training on a large-scale, highly heterogeneous collection of client hardware, increased communication costs, and frequently slower convergence (Yousefpour et al., 2023).
@@
Quantifying and lowering the carbon emissions associated with machine learning (ML) training and inference in a data center environment has gained attention recently (Strubell et al., 2020; Lacoste et al., 2019; Henderson et al., 2020). Federated Learning’s (FL) carbon footprint and the elements that make FL carbon efficient have not yet been fully investigated. Previous studies have provided preliminary results, but they have only provided a partial picture by quantifying the carbon effects of FL in a simulation context or under numerous simplifying assumptions (Qiu et al., 2021; Wu et al., 2022b). These works refrained from exploring dimensions of the design space toward attaining Green FL, instead concentrating on measures and opportunity sizing. In (Yousefpour et al., 2023), the authors analyzed that a strong correlation exists among all factors and artifacts between an FL task’s carbon footprint and the sum of its operating time and the number of users engaged in training or concurrency. Moreover, the authors concluded that the number of participating clients (concurrency) and training time to achieve the desired accuracy are directly related to the carbon emission of any FL framework. On the other hand, other factors like learning rates, batch sizes, aggregation targets, and local epochs affect how quickly the FL model converges to the target accuracy, affecting how long it takes to finish the training process. Although these factors have no direct impact on carbon emissions, they do indirectly affect the total training speed. Therefore, these parameters should be included in the multidimensional design of Green FL’s “time” and “performance” aspects. Conversely, concurrency directly impacts time and carbon emissions and hence needs to be taken into account more while designing FL systems to lower carbon emissions.
@@
The concept of Green Federated Learning was also studied by Yousefpour et al. (Yousefpour et al., 2023). In this study, the authors introduced the notion of Green FL, which entails FL parameter optimization and design decisions to reduce carbon emissions while maintaining competitive performance and training efficiency. Two contributions were made to this work. First, a data-driven methodology to directly measure FL jobs that are executed on a scale in the real world on millions of phones to quantify the carbon emissions of FL. Second, a discussion on difficulties, recommendations, and takeaways from researching the trade-off between energy efficiency, performance, and training time in an FL system. Their findings lay a solid platform for future research in the field of green AI and provide useful insights into how FL can reduce its carbon impact.
6.1. Insight on assessing Energy and Carbon Footptint
@@
The SOTA surveys related to FL frameworks are divided into various challenges: privacy, heterogeneity, client selection, model compression, and communications rounds. Standard accuracy metrics, such as the handling of heterogeneous data, reliability, response time, convergence rate, number of communication rounds, and accuracy of a given model, usually determine SOTA results. However, Energy efficiency or emissions of carbon footprint are frequently disregarded. While the explored metrics are still important, we want to inspire academics to pay attention to other indicators that correspond with the growing concern of the civil community about global warming. Hence, we survey the works where the inclusion of released CO2 is considered a critical parameter in calculating carbon emissions for FL by quantifying it. Qiu et al.(Qiu et al., 2021) presented the first comprehensive analysis of FL’s carbon footprint. The authors proposed a strict model to measure the carbon footprint, making it easier to look into the connection between FL design and carbon emissions. In this work, the authors considered two main procedures to calculate the environmental cost of training deep learning models on edge or in data centers. First, the authors analyzed the energy used by the approach, primarily considering the overall energy used by the hardware. The latter amount is then translated, using geographic coordinates, into CO2 emissions. On the other hand, Yang et al.(Yang et al., 2021) examined the issue of FL across wireless communication networks: energy-efficient transmission and computing resource allocation. The model under consideration involves the utilization of restricted local computational resources by individual users to train a local FL model using the data they have collected. Subsequently, the trained FL model is transmitted to a federated server (FS), which combines the local FL model and disseminates it to all users. The learning accuracy level affects both computation and communication latencies in FL since it includes users and the BS exchanging a learning model. Meanwhile, both local computing energy and transmission energy must be taken into account during the FL process due to the wireless consumers’ limited energy budget. Based on this concept, Qiu et al.(Qiu et al., 2023) studied the energy needed by the procedure, primarily considering the overall energy used by the hardware. It consists of communication and training energy for FL and centralized learning. The latter figure is then translated into CO2 emissions according to geographic areas, which will differ significantly depending on the energy sources. Since information on emissions associated with hardware manufacturing is still primarily unavailable, it is not included in this study. Savazzi et al.(Savazzi et al., 2023) provided a revolutionary framework for FL that allows for the analysis of energy and carbon footprints. The methodology under consideration measures the energy footprints and carbon equivalent emissions of both fully decentralized, consensus-based systems and vanilla FL solutions. The authors evaluated the estimated carbon and energy footprints while taking into account the effects of different communication environments and the size of the learner community. The suggested framework sheds light on how the various elements of the FL designs affect carbon emissions and energy consumption in terms of system accuracy and number of rounds.
@@
Reference Proposal Evaluation Criteria Results Communication Rounds Client Selection Model Compression FedGreen (Li et al., 2021b) Create a fine-grained gradient compression technique by fusing server-side element-wise aggregation with device-side gradient reduction X X Give a trade-off analysis between learning accuracy and energy efficiency for FL using gradient compression. GREED (Albelaihi et al., 2022) Propose a client selection mechanism to reduce the energy consumption for IoT devices X X Ascertain that every client who is chosen has enough energy to upload their local models before the deadline by finding the best balance between increasing the number of clients that are chosen and reducing the energy that is taken from their batteries. (Thakur et al., 2023b) Propose a client selection algorithm for activity recognition with smartphone X X chooses the greatest number of clients in order to improve FL’s model correctness and convergence rate. (Salh et al., 2023) create a productive way to integrate joint edge intelligence nodes in order to reduce energy usage. X X By jointly optimizing the CPU frequency, improve calculation and transmission to meet FL time and minimize energy usage for IoT devices.
6.2. Application of Green Federated Learning
@@
Little evidence of energy-efficient FL (green FL) for IoT applications is available in the literature. A few recent studies have proposed integrating FL into IoT systems to provide a sustainable environment, as summarized in Table 6 as follows: Li et al.(Li et al., 2021b) suggested FedGreen, which adds fine-grained gradient compression to the original FL to improve it and control the devices’ overall energy consumption. The authors introduced the necessary operations to make gradient compression in FL easier, such as server-side element-wise aggregation and device-side gradient reduction. The compressed local gradients’ contributions to various compression ratios were assessed using data from a publicly available dataset. Additionally, the authors looked at a trade-off between learning precision and energy efficiency and derived the ideal compression ratio and processing frequency for each device. In order to maximize the number of selected clients and minimize the energy taken from batteries for the selected clients, Rana et al.(Albelaihi et al., 2022) proposed a novel client selection called “EnerGy-AwaRe CliEnt SElection for Green FeDerated Learning (GREED)” that ensures that all selected clients have enough energy to upload their local models before the deadline. Thakur et al.(Thakur et al., 2023b) proposed an energy-aware FL framework for activity recognition using a simple client selection algorithm based on the idea that participation of more clients reduces the number of communications rounds. Hence, it reduces energy consumption. Salh et al.(Salh et al., 2023) proposed an energy-efficient FL framework for green IoT to minimize energy consumption and meet the FL time requirement for all IoT devices. It does this by looking into energy-efficient bandwidth allocation, computing “Central Processing Unit” (CPU) frequency, optimization transmission power, and the desired level of learning accuracy. By resolving the bandwidth optimization problem in closed form, the approach effectively optimized the compute frequency allocation and decreased energy consumption in IoT devices. An “Alternative Direction Algorithm” (ADA) could resolve the remaining computational frequency allocation, transmission power allocation, and loss to lower energy consumption and complexity at each FL time iteration from IoT devices to edge intelligence nodes.
7. FL based Green- Aware IoT Applications
@@
(Wang et al., 2022) Presented a federated transfer learning framework for cross-domain in smart manufacturing X Boost accuracy and response speed in comparison to more complex techniques
@@
(Yazdinejad et al., 2022) Create a threat hunting system called block hunter using federated learning to automatically search for attacks in IIoT networks based on blockchain technology for smart factories to prevent cyberattacks. Improve the accuracy of detection and the speed of response.
@@
(Huong et al., 2022) Proposed an interpretable FedX architecture to identify unknown attacks in smart IoT-based factories Enhance accuracy with faster training time & less memory consumption
@@
(Fan et al., 2023) provide a privacy-preserving data aggregation approach for the IIoT based on federated learning (FLPDA). Data aggregation to defend against reverse analysis threats from industry administration centers and preserve individual user model modifications in federated learning. X Compared to current systems, it has less computational, storage, and communication overhead with enhanced security & privacy.
@@
(Savazzi et al., 2023) Create an analytical framework for energy and carbon footprints for environmentally friendly and sustainable FL-based IIoT systems. A trade-off between energy and test accuracy should be made in light of the model and data footprints for the intended industrial applications.
@@
(Peres et al., 2023) Propose conceptual FL framework for collaborative industrial AI applications Improve security, privacy, and execution speed while lowering computer costs.
@@
(Salh et al., 2023) In order to minimize energy consumption and meet the FL time requirement for all IoT devices, develop an efficient integration of joint edge intelligence nodes. It does this by looking into energy-efficient bandwidth allocation, computing Central Processing Unit (CPU) frequency, optimization transmission power, and the desired level of learning accuracy. By resolving the bandwidth optimization problem in closed form, the approach effectively optimized the compute frequency allocation and decreased energy consumption in IoT devices. An Alternative Direction Algorithm (ADA) could be used to resolve the remaining computational frequency allocation, transmission power allocation, and loss in order to lower energy consumption and complexity at each FL time iteration from IoT devices to edge intelligence nodes.
@@
Massive amounts of data are constantly being generated by the growing number of Internet of Things (IoT) devices, but the current cloud-centric approach to IoT big data analysis has sparked worries among the public about data privacy and network costs (Xu et al., 2021). In current deployments, a cloud server centrally hosts and processes the real-time data that IoT devices collect. However, the sharing of privacy-sensitive data via cloud-based model training raises concerns from the public and may result in unacceptably high traffic loads and latency (Lim et al., 2020b). It is better to separate the requirement for remotely gathering and centrally processing the raw data from the model training process in order to lessen these difficulties. Given these drawbacks of the cloud-centric solution, scientists have recently introduced Federated Learning (FL), a novel method for training a shared global model on datasets decentralized at a loose federation of participating devices (clients) (McMahan et al., 2016).
@@
However, the training overhead presents FL with a significant hurdle. Popular DNN models, on the one hand, typically have tens of thousands to hundreds of millions of parameters (Sze et al., 2017). The training of the model is highly computationally and memory-intensive due to the constantly growing network complexity and training data (Lim et al., 2020b) resulting in a large carbon footprint. However, frequent and high-dimensional parameter changes come with a high cost of communication and can lead to a training bottleneck (Tao and Li, 2018). perhaps worse, because IoT devices typically have limited computational power, meomory and transmission capacity, the training period will be excessively lengthy or perhaps intolerable. This makes it difficult to quickly implement and use AI models in real-world settings. In this section, we survey the application of green-aware AI based on federated learning for different IoT applications such as smart industry, smart cities, smart healthcare, UAVs, and intelligent transportation.
@@
(Otoum et al., 2022) In order to achieve network security and trustworthiness in smart cities, present an adaptive framework that combines blockchain technology with federated learning. X Demonstrate the efficacy of the framework in terms of network longevity, energy consumption, and trust using different parameters. The suggested model continues to have high accuracy and detection rates.
(Zhang et al., 2022) Propose an FL framework to overcome communication overhead issues in the application with smart cities. X Demonstrate exceptional training efficiency, precise prediction accuracy, and resilience to erratic network settings.
(Algarni et al., 2023) Provide a FL-based privacy-aware plan for private remote sensing information. Increase detection rate, accuracy, and convergence speed over current solutions.
(El Houda et al., 2023) Propose DETECT, a promising edge-based architecture that uses FL to secure Internet of Things applications. By protecting the privacy of MEC collaborators and, by extension, the privacy of IoT devices, DETECT enables several MEC domains to work together to jointly and securely reduce IoT assaults. X Utilizing the Edge-IIoTset and NSL-KDD datasets, two well-known IoT attack scenarios, DETECT demonstrates a noteworthy accuracy in terms of both F1 score (87% in NSL-KDD and 99% in Edge-IIoTset) and Accuracy (86% in NSL-KDD and 99% in Edge-IIoTset).
7.1. Smart Industry
@@
(Akter et al., 2022) Create a secure and private FL-based IoMT system. X After specific iterations, the findings showed a 90% accuracy performance, superior to other baseline approaches with accuracy levels of about 80% using the same epsilon number of 4. Furthermore, this architecture satisfies the privacy preservation paradigm more successfully and quickly.
@@
(Qayyum et al., 2022) Employed the recently developed idea of clustered FL to perform an automated diagnostic of COVID-19. X Increase forecast accuracy while cutting down on time expenditure.
@@
(Guo et al., 2022a) Provide an FL-based real-time IoMT system for medical data processing. X X Faster service response times for IoMT systems operating in real-time.
@@
(Wu et al., 2022a) Provide an innovative cloud-edge FL architecture for in-home health monitoring that protects user privacy by storing user data locally and learning a shared global model in the cloud from numerous houses at the network edges. Derive precise and individualized health tracking by fine-tuning the model using a dataset that is generated from the user’s personal, class-balanced data with reduced communication cost.
@@
(Zhang et al., 2023b) Provide an end-to-end federated representation learning approach for human mobility predictions To tackle the constraints of heterogeneity in mobility patterns and data scarcity. Advantageous over SOTA methods in terms of security, accuracy and response time.
@@
(Gupta et al., 2023b) Develop a federated edge-based framework for IoMT devices with limited resources. X Reduce training time of videos by 3.6 hrs when employing the split learning phase. Also, achieve 90.32% global accuracy at the server.
@@
(Rehman et al., 2023) Develop a framework for attack detection based on FL. X X Increase the attack detection rate accuracy to 80.09%.
@@
(Kalapaaking et al., 2023) Provide a blockchain-based federated learning solution that protects healthcare systems from poisoning threats by utilizing SMPC model verification. X X Boost security and privacy for intelligent IoMT applications.
@@
(Wang et al., 2023b) Suggest a Privacy Protection Framework for Edge Computing Federated Learning (PPFLEC) IoMT applications X Reduce time expenditure by 40% in comparison to current methods with security measure.
@@
(Yu et al., 2023) In order to overcome the idea drift and convergence instability difficulties in the online federated learning process to recognize physical activities, an unsupervised gradient aggregation technique is created along with an algorithm to calculate unsupervised gradients under the consistency training proposition. X Achieve an average additional 10% improvement in accuracy and response time.
7.2. Smart Cities
@@
The installation of 5G base stations, intercity high-speed rail and urban rail transit, artificial intelligence, and industrial Internet have all an impact on energy conservation and the reduction of CO2 emissions in smart cities (Guo et al., 2022b). Smart cities are actually the result of combining various infrastructure, technologies, and solutions to create intelligent metropolitan areas. Real-time smart IoT systems are one of the main aspects of smart cities. Combining FL and IoT applications is suggested as a means of achieving this, and Table 8 shows some encouraging outcomes.

### v2

@@ Introduction
In the previous section Challenges and solutions in Federated Learning, several FL-related challenges and solutions are discussed elaborately. However, how FL can affect the ecosystem is still unknown and unclear. There is a relation between FL design and carbon emission [Qiu et al.(2021)]. According to the survey, an energy-efficient FL framework reduces the carbon footprint. The potential for ML systems to greatly increase carbon emissions exists. Recent research [Schwartz et al.(2019), Strubell et al.(2020)] has recommended a number of mitigating techniques and used case studies to illustrate these potential implications. To more accurately assess the overall energy and carbon footprints of machine learning—in both research and production settings—systematic and precise measurements are required. Among other advantages, accurate accounting of carbon and energy impacts promotes awareness, motivates mitigation efforts, and aligns incentives with energy efficiency [Schwartz et al.(2019)]. However, energy or carbon emissions data are not routinely reported in the majority of ML research articles. Recent research primarily looks into how machine learning research affects the climate. By analyzing the anticipated power consumption and carbon emissions for a series of case studies, Strubell et al. [Strubell et al.(2020)] illustrated the problem of carbon and energy consequences of big NLP models. According to the authors, “researchers should prioritize computationally efficient hardware and algorithms”, “academic researchers need equitable access to computation resources”, and “authors should report training time and sensitivity to hyperparameters.” Similar recommendations are made by Schwartz et al. [Schwartz et al.(2019)], who propose floating point operations (FPOs) as a guiding efficiency parameter. A website for calculating carbon emissions based on GPU kind, experiment duration, and cloud provider was just made available by Lacoste et al. [Lacoste et al.(2019)]. Similarly, FL - a cooperative machine learning method that uses data from decentralized entities to train a centralized model—can also be resource-intensive and have a significant carbon impact, mainly when used on a large scale. In contrast to centralized artificial intelligence (AI), which may reliably utilize renewable energy sources at carefully positioned data centers, cross-device FL can utilize up to hundreds of millions of end-user devices spread throughout the globe that use a variety of energy sources. According to a recent study, training a model with FL can emit up to 80 kg of carbon dioxide equivalent (CO2e), which is more than developing a larger transformer model in a centralized training setup using AI accelerators [Wu et al.(2022b)]. Several reasons contribute to the relative inefficiency: the overhead of training on a large-scale, highly heterogeneous collection of client hardware, increased communication costs, and frequently slower convergence [Yousefpour et al.(2023)].
@@
Quantifying and lowering the carbon emissions associated with machine learning (ML) training and inference in a data center environment has gained attention recently [Strubell et al.(2020), Lacoste et al.(2019), Henderson et al.(2020)]. Federated Learning’s (FL) carbon footprint and the elements that make FL carbon efficient have not yet been fully investigated. Previous studies have provided preliminary results, but they have only provided a partial picture by quantifying the carbon effects of FL in a simulation context or under numerous simplifying assumptions [Qiu et al.(2021), Wu et al.(2022b)]. These works refrained from exploring dimensions of the design space toward attaining Green FL, instead concentrating on measures and opportunity sizing. In [Yousefpour et al.(2023)], the authors analyzed that a strong correlation exists among all factors and artifacts between an FL task’s carbon footprint and the sum of its operating time and the number of users engaged in training or concurrency. Moreover, the authors concluded that the number of participating clients (concurrency) and training time to achieve the desired accuracy are directly related to the carbon emission of any FL framework. On the other hand, other factors like learning rates, batch sizes, aggregation targets, and local epochs affect how quickly the FL model converges to the target accuracy, affecting how long it takes to finish the training process. Although these factors have no direct impact on carbon emissions, they do indirectly affect the total training speed. Therefore, these parameters should be included in the multidimensional design of Green FL’s “time” and “performance” aspects. Conversely, concurrency directly impacts time and carbon emissions and hence needs to be taken into account more while designing FL systems to lower carbon emissions.
@@
The concept of Green Federated Learning was also studied by Yousefpour et al. [Yousefpour et al.(2023)]. In this study, the authors introduced the notion of Green FL, which entails FL parameter optimization and design decisions to reduce carbon emissions while maintaining competitive performance and training efficiency. Two contributions were made to this work. First, a data-driven methodology to directly measure FL jobs that are executed on a scale in the real world on millions of phones to quantify the carbon emissions of FL. Second, a discussion on difficulties, recommendations, and takeaways from researching the trade-off between energy efficiency, performance, and training time in an FL system. Their findings lay a solid platform for future research in the field of green AI and provide useful insights into how FL can reduce its carbon impact.
Insight on assessing Energy and Carbon Footptint
@@
The SOTA surveys related to FL frameworks are divided into various challenges: privacy, heterogeneity, client selection, model compression, and communications rounds. Standard accuracy metrics, such as the handling of heterogeneous data, reliability, response time, convergence rate, number of communication rounds, and accuracy of a given model, usually determine SOTA results. However, Energy efficiency or emissions of carbon footprint are frequently disregarded. While the explored metrics are still important, we want to inspire academics to pay attention to other indicators that correspond with the growing concern of the civil community about global warming. Hence, we survey the works where the inclusion of released CO2 is considered a critical parameter in calculating carbon emissions for FL by quantifying it. Qiu et al.[Qiu et al.(2021)] presented the first comprehensive analysis of FL’s carbon footprint. The authors proposed a strict model to measure the carbon footprint, making it easier to look into the connection between FL design and carbon emissions. In this work, the authors considered two main procedures to calculate the environmental cost of training deep learning models on edge or in data centers. First, the authors analyzed the energy used by the approach, primarily considering the overall energy used by the hardware. The latter amount is then translated, using geographic coordinates, into CO2 emissions. On the other hand, Yang et al.[Yang et al.(2021)] examined the issue of FL across wireless communication networks: energy-efficient transmission and computing resource allocation. The model under consideration involves the utilization of restricted local computational resources by individual users to train a local FL model using the data they have collected. Subsequently, the trained FL model is transmitted to a federated server (FS), which combines the local FL model and disseminates it to all users. The learning accuracy level affects both computation and communication latencies in FL since it includes users and the BS exchanging a learning model. Meanwhile, both local computing energy and transmission energy must be taken into account during the FL process due to the wireless consumers’ limited energy budget. Based on this concept, Qiu et al.[Qiu et al.(2023)] studied the energy needed by the procedure, primarily considering the overall energy used by the hardware. It consists of communication and training energy for FL and centralized learning. The latter figure is then translated into CO2 emissions according to geographic areas, which will differ significantly depending on the energy sources. Since information on emissions associated with hardware manufacturing is still primarily unavailable, it is not included in this study. Savazzi et al.[Savazzi et al.(2023)] provided a revolutionary framework for FL that allows for the analysis of energy and carbon footprints. The methodology under consideration measures the energy footprints and carbon equivalent emissions of both fully decentralized, consensus-based systems and vanilla FL solutions. The authors evaluated the estimated carbon and energy footprints while taking into account the effects of different communication environments and the size of the learner community. The suggested framework sheds light on how the various elements of the FL designs affect carbon emissions and energy consumption in terms of system accuracy and number of rounds.
@@
Reference Proposal Evaluation Criteria Results Communication Rounds Client Selection Model Compression FedGreen [Li et al.(2021b)] Create a fine-grained gradient compression technique by fusing server-side element-wise aggregation with device-side gradient reduction X X Give a trade-off analysis between learning accuracy and energy efficiency for FL using gradient compression. GREED [Albelaihi et al.(2022)] Propose a client selection mechanism to reduce the energy consumption for IoT devices X X Ascertain that every client who is chosen has enough energy to upload their local models before the deadline by finding the best balance between increasing the number of clients that are chosen and reducing the energy that is taken from their batteries. [Thakur et al.(2023b)] Propose a client selection algorithm for activity recognition with smartphone X X chooses the greatest number of clients in order to improve FL’s model correctness and convergence rate. [Salh et al.(2023)] create a productive way to integrate joint edge intelligence nodes in order to reduce energy usage. X X By jointly optimizing the CPU frequency, improve calculation and transmission to meet FL time and minimize energy usage for IoT devices.
Application of Green Federated Learning
@@
Little evidence of energy-efficient FL (green FL) for IoT applications is available in the literature. A few recent studies have proposed integrating FL into IoT systems to provide a sustainable environment, as summarized in Table 6 as follows: Li et al.[Li et al.(2021b)] suggested FedGreen, which adds fine-grained gradient compression to the original FL to improve it and control the devices’ overall energy consumption. The authors introduced the necessary operations to make gradient compression in FL easier, such as server-side element-wise aggregation and device-side gradient reduction. The compressed local gradients’ contributions to various compression ratios were assessed using data from a publicly available dataset. Additionally, the authors looked at a trade-off between learning precision and energy efficiency and derived the ideal compression ratio and processing frequency for each device. In order to maximize the number of selected clients and minimize the energy taken from batteries for the selected clients, Rana et al.[Albelaihi et al.(2022)] proposed a novel client selection called “EnerGy-AwaRe CliEnt SElection for Green FeDerated Learning (GREED)” that ensures that all selected clients have enough energy to upload their local models before the deadline. Thakur et al.[Thakur et al.(2023b)] proposed an energy-aware FL framework for activity recognition using a simple client selection algorithm based on the idea that participation of more clients reduces the number of communications rounds. Hence, it reduces energy consumption. Salh et al.[Salh et al.(2023)] proposed an energy-efficient FL framework for green IoT to minimize energy consumption and meet the FL time requirement for all IoT devices. It does this by looking into energy-efficient bandwidth allocation, computing “Central Processing Unit” (CPU) frequency, optimization transmission power, and the desired level of learning accuracy. By resolving the bandwidth optimization problem in closed form, the approach effectively optimized the compute frequency allocation and decreased energy consumption in IoT devices. An “Alternative Direction Algorithm” (ADA) could resolve the remaining computational frequency allocation, transmission power allocation, and loss to lower energy consumption and complexity at each FL time iteration from IoT devices to edge intelligence nodes.
FL based Green- Aware IoT Applications
@@
[Wang et al.(2022)] Presented a federated transfer learning framework for cross-domain in smart manufacturing X Boost accuracy and response speed in comparison to more complex techniques
@@
[Yazdinejad et al.(2022)] Create a threat hunting system called block hunter using federated learning to automatically search for attacks in IIoT networks based on blockchain technology for smart factories to prevent cyberattacks. Improve the accuracy of detection and the speed of response.
@@
[Huong et al.(2022)] Proposed an interpretable FedX architecture to identify unknown attacks in smart IoT-based factories Enhance accuracy with faster training time & less memory consumption
@@
[Fan et al.(2023)] provide a privacy-preserving data aggregation approach for the IIoT based on federated learning (FLPDA). Data aggregation to defend against reverse analysis threats from industry administration centers and preserve individual user model modifications in federated learning. X Compared to current systems, it has less computational, storage, and communication overhead with enhanced security & privacy.
@@
[Savazzi et al.(2023)] Create an analytical framework for energy and carbon footprints for environmentally friendly and sustainable FL-based IIoT systems. A trade-off between energy and test accuracy should be made in light of the model and data footprints for the intended industrial applications.
@@
[Peres et al.(2023)] Propose conceptual FL framework for collaborative industrial AI applications Improve security, privacy, and execution speed while lowering computer costs.
@@
[Salh et al.(2023)] In order to minimize energy consumption and meet the FL time requirement for all IoT devices, develop an efficient integration of joint edge intelligence nodes. It does this by looking into energy-efficient bandwidth allocation, computing Central Processing Unit (CPU) frequency, optimization transmission power, and the desired level of learning accuracy. By resolving the bandwidth optimization problem in closed form, the approach effectively optimized the compute frequency allocation and decreased energy consumption in IoT devices. An Alternative Direction Algorithm (ADA) could be used to resolve the remaining computational frequency allocation, transmission power allocation, and loss in order to lower energy consumption and complexity at each FL time iteration from IoT devices to edge intelligence nodes.
@@
Massive amounts of data are constantly being generated by the growing number of Internet of Things (IoT) devices, but the current cloud-centric approach to IoT big data analysis has sparked worries among the public about data privacy and network costs [Xu et al.(2021)]. In current deployments, a cloud server centrally hosts and processes the real-time data that IoT devices collect. However, the sharing of privacy-sensitive data via cloud-based model training raises concerns from the public and may result in unacceptably high traffic loads and latency [Lim et al.(2020b)]. It is better to separate the requirement for remotely gathering and centrally processing the raw data from the model training process in order to lessen these difficulties. Given these drawbacks of the cloud-centric solution, scientists have recently introduced Federated Learning (FL), a novel method for training a shared global model on datasets decentralized at a loose federation of participating devices (clients) [McMahan et al.(2016)].
@@
However, the training overhead presents FL with a significant hurdle. Popular DNN models, on the one hand, typically have tens of thousands to hundreds of millions of parameters [Sze et al.(2017)]. The training of the model is highly computationally and memory-intensive due to the constantly growing network complexity and training data [Lim et al.(2020b)] resulting in a large carbon footprint. However, frequent and high-dimensional parameter changes come with a high cost of communication and can lead to a training bottleneck [Tao and Li(2018)]. perhaps worse, because IoT devices typically have limited computational power, meomory and transmission capacity, the training period will be excessively lengthy or perhaps intolerable. This makes it difficult to quickly implement and use AI models in real-world settings. In this section, we survey the application of green-aware AI based on federated learning for different IoT applications such as smart industry, smart cities, smart healthcare, UAVs, and intelligent transportation.
@@
[Otoum et al.(2022)] In order to achieve network security and trustworthiness in smart cities, present an adaptive framework that combines blockchain technology with federated learning. X Demonstrate the efficacy of the framework in terms of network longevity, energy consumption, and trust using different parameters. The suggested model continues to have high accuracy and detection rates.
[Zhang et al.(2022)] Propose an FL framework to overcome communication overhead issues in the application with smart cities. X Demonstrate exceptional training efficiency, precise prediction accuracy, and resilience to erratic network settings.
[Algarni et al.(2023)] Provide a FL-based privacy-aware plan for private remote sensing information. Increase detection rate, accuracy, and convergence speed over current solutions.
[El Houda et al.(2023)] Propose DETECT, a promising edge-based architecture that uses FL to secure Internet of Things applications. By protecting the privacy of MEC collaborators and, by extension, the privacy of IoT devices, DETECT enables several MEC domains to work together to jointly and securely reduce IoT assaults. X Utilizing the Edge-IIoTset and NSL-KDD datasets, two well-known IoT attack scenarios, DETECT demonstrates a noteworthy accuracy in terms of both F1 score (87% in NSL-KDD and 99% in Edge-IIoTset) and Accuracy (86% in NSL-KDD and 99% in Edge-IIoTset).
Smart Industry
@@
[Akter et al.(2022)] Create a secure and private FL-based IoMT system. X After specific iterations, the findings showed a 90% accuracy performance, superior to other baseline approaches with accuracy levels of about 80% using the same epsilon number of 4. Furthermore, this architecture satisfies the privacy preservation paradigm more successfully and quickly.
@@
[Qayyum et al.(2022)] Employed the recently developed idea of clustered FL to perform an automated diagnostic of COVID-19. X Increase forecast accuracy while cutting down on time expenditure.
@@
[Guo et al.(2022a)] Provide an FL-based real-time IoMT system for medical data processing. X X Faster service response times for IoMT systems operating in real-time.
@@
[Wu et al.(2022a)] Provide an innovative cloud-edge FL architecture for in-home health monitoring that protects user privacy by storing user data locally and learning a shared global model in the cloud from numerous houses at the network edges. Derive precise and individualized health tracking by fine-tuning the model using a dataset that is generated from the user’s personal, class-balanced data with reduced communication cost.
@@
[Zhang et al.(2023b)] Provide an end-to-end federated representation learning approach for human mobility predictions To tackle the constraints of heterogeneity in mobility patterns and data scarcity. Advantageous over SOTA methods in terms of security, accuracy and response time.
@@
[Gupta et al.(2023b)] Develop a federated edge-based framework for IoMT devices with limited resources. X Reduce training time of videos by 3.6 hrs when employing the split learning phase. Also, achieve 90.32% global accuracy at the server.
@@
[Rehman et al.(2023)] Develop a framework for attack detection based on FL. X X Increase the attack detection rate accuracy to 80.09%.
@@
[Kalapaaking et al.(2023)] Provide a blockchain-based federated learning solution that protects healthcare systems from poisoning threats by utilizing SMPC model verification. X X Boost security and privacy for intelligent IoMT applications.
@@
[Wang et al.(2023b)] Suggest a Privacy Protection Framework for Edge Computing Federated Learning (PPFLEC) IoMT applications X Reduce time expenditure by 40% in comparison to current methods with security measure.
@@
[Yu et al.(2023)] In order to overcome the idea drift and convergence instability difficulties in the online federated learning process to recognize physical activities, an unsupervised gradient aggregation technique is created along with an algorithm to calculate unsupervised gradients under the consistency training proposition. X Achieve an average additional 10% improvement in accuracy and response time.
Smart Cities
@@
The installation of 5G base stations, intercity high-speed rail and urban rail transit, artificial intelligence, and industrial Internet have all an impact on energy conservation and the reduction of CO2 emissions in smart cities [Guo et al.(2022b)]. Smart cities are actually the result of combining various infrastructure, technologies, and solutions to create intelligent metropolitan areas. Real-time smart IoT systems are one of the main aspects of smart cities. Combining FL and IoT applications is suggested as a means of achieving this, and Table 8 shows some encouraging outcomes.